If someone runs 3 3/5 miles in 33 mins. What is his pace in miles per hour

Answer by amarjeeth123(574)

Distance covered=3 3/5 miles=3.6 miles
Time taken=33 minutes=33/60 hours=0.55 hours
Speed=Distance covered/Time taken=3.6/0.55=6.54 miles/hour